The heart consumes 350 kJ of energy per day. What is heart efficiency if it strikes 62 beats per minute and makes a work of 1.1 J per one stroke?


1
Expert's answer
2021-10-11T16:45:07-0400

The efficiency is given as follows:


"\eta = \dfrac{W_{done}}{Q_{cons}}\times 100\%"

where "W_{done}" is the useful work done by the heart, "Q_{cons} = 350kJ" is the consumed energy. The useful work can be found as follows:


"W_{done} = 1.1J\times 62\times 1440 = 98208J = 98.208kJ"

where "1440" is the number of minutes in one day.

Thus, obtain:


"\eta = \dfrac{98.208kJ}{350kJ}\times 100\% \approx 28\%"

Answer. 28%.